OpenStudy (anonymous):

find the simultaneous solution y+3x+11y=99 5-15x+21y=77

OpenStudy (perl):

you can simplify the equations 12y + 3x = 99 -15x + 21y = 72

OpenStudy (anonymous):

once they are simplified what then

OpenStudy (perl):

you can solve by elimination or substitution

OpenStudy (anonymous):

can u demonstrate?

OpenStudy (anonymous):

i have tried elimination but its not working out